Define adiabatic lapse rate and show that
dT/dh= {(–γ–1)/γ} (Mg/R)
Plot a graph between atmospheric temperature with altitude.
1
Expert's answer
2020-02-27T10:12:19-0500

The air parcel undergoes only adiabatic transformations "dQ=0", and the atmosphere is in hydrostatic equilibrium.


"dp=-g\cdot \rho \cdot dh"


"\frac{dp}{\rho}=-g\cdot dh" (1)


"-dA=dU\to -pdV=mc_VdT"


"d(pV)=pdV+Vdp"


"Vdp-d(pV)=mc_VdT"


From the ideal gas low "pV=nRT"


"Vdp-d(nRT)=Vdp-nRdT=mc_VdT"

"\frac{1}{\rho}dp-\frac{n}{nM}RdT=c_VdT\to \frac{1}{\rho}dp=(\frac{R}{M}+c_V)dT\to"


"\frac {dp}{\rho}=c_pdT" (2)


"c_pdT=-g\cdot dh"


Combining (1) with the equation (2), yields


"\frac{dT}{dh}=-\frac{g}{c_p}" or "\frac{dT}{dh}=-\frac{g}{\frac{\gamma R}{(\gamma-1)M}}\to \frac{dT}{dh}=-\frac{\gamma-1}{\gamma}\frac{Mg}{R}"